# Unsupervised learning in particle physics

### v1

@@
Unsupervised learning in particle physics
[…]
${}^{a}$ Center for Computational Natural Sciences and Bioinformatics, International Institute of Information Technology, Hyderabad 500 032, India
${}^{b}$ Indian Institute of Science Education and Research Thiruvananthapuram, Vithura, Kerala, 695 551, India
${}^{c}$ Center for Quantum Science and Technology, International Institute of Information Technology, Hyderabad 500 032, India
${}^{d}$ iHub-Data, International Institute of Information Technology, Hyderabad 500 032, India
We review the main applications of unsupervised learning methods in particle physics, i.e., clustering, anomaly detection, detector simulation, and unfolding. Unsupervised methods are ideal for anomaly detection tasks—machine learning models can be trained on background data to identify deviations if we model the background data precisely. The learning can also be partially unsupervised when we can provide some information about the anomalies at the data level. Generative models are useful in speeding up detector simulations—they can mimic the computationally intensive task without large resources. They can also efficiently map detector-level data to parton-level data (i.e., data unfolding). In this review, we focus on interesting ideas and connections and briefly overview the underlying techniques wherever necessary.
jai.bardhan90@googlemail.com
tanumoy@iisertvm.ac.in
subhadip.mitra@iiit.ac.in
cyrin.neeraj@research.iiit.ac.in
monalisa.p@ihub-data.iiit.ac.in
1 Introduction
In recent years—especially since the advent of deep learning—machine learning (ML) techniques have undergone what can be termed dramatic revolutions. Their performance and range of applicability have increased manyfold. ML tools are now standard in almost every field of study. Particle physics is no exception. These techniques are not new to particle physics—from the LEP to the LHC, high-energy data have been analysed with multivariate techniques like boosted decision trees, neural networks (NNs), etc. (see the references in, e.g., Ref. [1]). The famous discovery of Higgs, which completed the experimental verification of the particle spectrum of the Standard Model (SM), relied on ML techniques. Recently, however, there has been a surge in interest in these methods. Modern and sophisticated ML techniques are being adapted to particle physics problems at an unprecedented rate (see Ref. [2] [^2] The Particle Data Book  [3] now includes a review on ML.) for various reasons. For example, precise measurements in some unexplored parts of the SM, like the Higgs self-coupling or light-quark Yukawa couplings, will be beyond the reach of the traditional analyses [4, 5, 6, 7]. Then, the fact that the LHC is yet to find any signature of physics beyond the SM (BSM) implies that we need to look for processes with even smaller cross-sections or leading to minor departures from known distributions. With the planned high luminosity upgrade (HL-LHC), the LHC will gather, analyse, and record a vast amount of collision data. Modern ML techniques are ideal for such tasks—they can handle tiny signals much better than the usual cut-based analyses designed by physicists’ intuitions or find (learn) patterns from a vast amount of data.
@@ 1 Introduction
Unsupervised tasks are not new to particle physics. Common jet clustering is essentially an unsupervised task as the algorithm (like anti-$k_{T}$ or Cambridge/Aachen) groups objects without any specific predefined target jet. [^3] In practice, some tuning of the parameters may be necessary to tag boosted objects; for example, see Refs. [8, 9], where a dijet signal from a boosted BSM singlet boson is tagged. We look at clustering tasks in Sec. 2. In Sec. 3, we focus on anomaly detection, another major task for nonsupervised methods. Anomaly detection without supervision is essentially finding deviations or departures (anomalies) from known patterns without any prior expectation about the nature of the anomalies (like knowledge of any NP model or interaction). If the ML model learns the distribution of the background (SM) events directly from the data, it can identify events not coming from that distribution (i.e., out-of-distribution events). In this section, we discuss data-driven background estimation, autoencoders, variational autoencoders, weakly supervised methods, topic modelling in jet space, and self-supervised learning methods. In the next section (Sec. 4), we look at how ML models, especially generative adversarial networks (GANs), can play a key role in fast detector simulations and unfolding the data distributions recorded at the detectors. Finally, we conclude in Sec. 5. One comment is in order before we move on to the next section. Most developments we discuss here are recent, and applications of nonsupervised techniques in particle physics are still nascent. [^4] The HEP community has put out several challenges related to unsupervised learning in the last five years: Fast Calorimeter Challenge: https://calochallenge.github.io/homepage/, Anomaly Detection Challenge: https://mpp-hep.github.io/ADC2021/, etc. Hence, in this review, we focus more on the methods rather than quantitative comparisons of the results. For specialised reviews on anomaly detection and generative modelling in detector simulations, see Ref. [10] and Ref. [11], respectively.
2 Clustering
@@
$\mathcal{T}$$\equiv$ $\{\bm{x}_{i}\}_{i=1}^{N_{T}}\stackrel{{\scriptstyle\textrm{iid}}}{{\sim}}\rho_{T}\,,$ (4)
$\mathcal{B}$$\equiv$ $\{\bm{x}^{\prime}_{i}\}_{i=1}^{N_{B}}\stackrel{{\scriptstyle\textrm{iid}}}{{\sim}}\rho_{B}\,.$ (5)
with $\{\bm{x}_{i}|\bm{x}_{i}\in\mathbb{R}^{D}\}_{i=1}^{N_{T}}$ and $\{\bm{x}^{\prime}_{i}|\bm{x}^{\prime}_{i}\in\mathbb{R}^{D}\}_{i=1}^{N_{B}}$ being two independent and identically distributed $D$-dimensional samples drawn independently from the probability density functions (PDFs) $\rho_{T}$ and $\rho_{B}$, respectively. The $\mathcal{T}$, $\mathcal{B}$ samples consist of $N_{T}$, $N_{B}$ points, respectively. In conventional approaches, the NP signals are simulated in specific models and compared with the experimental data to test for those models. Since the theoretical landscape for NP signals is enormous, there is always a possibility that an actual signal may be missed. This method looks for deviations between the simulated SM background ($\mathcal{B}$ samples) and the true data ($\mathcal{T}$), agnostic to any NP scenarios. They check to what significance the two samples are compatible with each other (i.e., $\rho_{B}=\rho_{T}$) by performing a statistical test of the null hypothesis $\{H_{0}:\rho_{T}=\rho_{B}\}$ versus the alternative hypothesis $\{H_{1}:\rho_{T}\neq\rho_{B}\}$. The test statistic (TS) quantifies how much the sample data deviates from what is expected under the null hypothesis. The TS over the trial sample is defined as:
$\textrm{TS}(\mathcal{B},\mathcal{T})\equiv\log\hat{\lambda}^{1/|\mathcal{T}|}=\frac{1}{N_{T}}\sum_{j=1}^{N_{T}}\log\frac{\hat{\rho}_{T}(\bm{x}_{j})}{\hat{\rho}_{B}(\bm{x}_{j})}\,,$ (6)
@@
$\hat{\rho}_{B}(\bm{x}_{j})$ $=$ $\frac{\overline{K}}{N_{B}}\frac{1}{\omega_{D}r_{j,B}^{D}},\quad\quad\hat{\rho}_{T}(\bm{x}_{j})=\frac{\overline{K}}{N_{T}-1}\frac{1}{\omega_{D}r_{j,T}^{D}}\,,$ (7)
(for any $\bm{x}_{j}\in\mathcal{T}$) where $\omega_{D}=\pi^{D/2}/\Gamma(D/2+1)$ is the volume of the unit sphere in $\mathbb{R}^{D}$, and $\overline{K}$ is the number of nearest neighbours.
@@ 3 Anomaly detection
Figure 1: ABCD method: Figure adapted from Ref. [15]. Signal (background) points are coloured blue (light brown). The independence of $f$ and $g$ implies Pr$(f\geq f_{c}{\rm\leavevmode\nobreak\ and\leavevmode\nobreak\ }g\geq g_{c})={\rm Pr}(f\geq f_{c})\times{\rm Pr}(g\geq g_{c})$.
@@
Figure 3: Illustration of the AE architecture (taken from Ref. [17]). An AE comprises two neural networks – the encoder and the decoder. The encoder learns a transformation from the input $x$ to a compressed (bottleneck, shown in orange) latent space representation $z$, and the decoder learns a transformation from the latent space representation $z$ to reconstructed output $x^{\prime}$. The AE is trained such that the reconstructed output is ideally identical to the input ($x\approx x^{\prime}$). AEs learn a data-biased algorithm that is excellent at compressing and reconstructing data. The reconstruction quality of $x^{\prime}$ with respect to $x$ can be modelled as an anomaly score.
Figure 3: Illustration of the AE architecture (taken from Ref. [17]). An AE comprises two neural networks – the encoder and the decoder. The encoder learns a transformation from the input $x$ to a compressed (bottleneck, shown in orange) latent space representation $z$, and the decoder learns a transformation from the latent space representation $z$ to reconstructed output $x^{\prime}$. The AE is trained such that the reconstructed output is ideally identical to the input ($x\approx x^{\prime}$). AEs learn a data-biased algorithm that is excellent at compressing and reconstructing data. The reconstruction quality of $x^{\prime}$ with respect to $x$ can be modelled as an anomaly score.
Figure 4: The graph shows the difference between the reconstruction loss of the signal and the background, for the AE trained on the SM background. A cut on the loss can be used to screen for anomalies.
@@
There are other methods that can learn the background distributions directly from the data; autoencoders and their variations are examples of such models.
3.1 Autoencoders
AEs are generally used for dimensionality reduction and feature learning, but they can also be used as unsupervised classifiers in some scenarios. In its basic form, an AE has a symmetrical architecture and consists of an encoder and a decoder, as shown in Fig. 4. The encoder consists of one or more layers of neurons that map the input data ($x\in\mathbb{R}^{n}$) into a lower-dimensional representation (latent space, $y\in\mathbb{R}^{k}$ with $k<n$). Most salient features of the input data captured by this compressed representation can be implemented by the function $f_{\theta}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{k}$. The number of neurons in the encoder decreases with each layer until it reaches the bottleneck layer. The bottleneck layer (also known as the latent space or the coding layer) has the lowest number of neurons and represents the space with the compressed form of the input data. After that, the decoder reconstructs the input data from the compressed representation, $g_{\varphi}:\mathbb{R}^{k}\rightarrow\mathbb{R}^{n}$. The decoder consists of the same number of layers as the encoder, with the number of neurons in each decoder layer increasing until it reaches the output layer, which has the same number of neurons as the input layer. Each hidden layer in this network typically uses activation functions like the rectified linear unit (ReLU) to introduce non-linearity. The autoencoder training involves minimising a loss function [e.g., mean-squared error (MSE)] that measures the difference between the input data and the reconstructed output:
@@ 3.1 Autoencoders
During training, the network learns to map the input data to a lower-dimensional representation and later reconstructs the input from that representation. As mentioned earlier, AEs can be used to identify events inconsistent with the expected background distribution of the SM as a sign of physics beyond the SM. The signal events would result in larger $L$ (Eq. (13)) than usual, i.e., a signal is more likely to manifest in the tail of the $L$ distribution, as shown in Fig. 4.
Figure 5: The distribution of the anomaly score from the AE is presented for the data and the five benchmark models ($H^{+}$, $W_{\text{KK}}$, $Z^{\prime}$, $W^{\prime}$ and $Z^{\prime}$, respectively). The predictions from these BSM models describe the expected number of events from ${}^{-1}$ of data for heavy particle masses around 2 TeV. The vertical lines represent the start of the three anomaly regions (AR). The labels associated with the three ARs state the visible cross section for the BSM processes which yields the same number of events as observed in the ${}^{-1}$ dataset. For further details refer to Ref. [25].
@@
Hidden-valley models can have an additional strong (confining) gauge group under which all SM particles are singlets [30]. However, these hidden dynamics might form light (dark) hadrons, which could be produced at the LHC through some higher-dimensional operators or an intermediary like the dark photon that mixes with the ordinary photon or the Higgs [e.g., the Higgs decaying to dark hadrons], etc. Like the bound states in QCD, dark hadrons can produce dark showers. Since modelling the signals is difficult in these cases, they make a good case for unsupervised learning in general. Ref. [31] demonstrated this by looking for soft-unclustered energy patterns (SUEPs) produced by dark showers-lead isotropic and soft energy deposits. The authors analysed one of the experimentally challenging scenarios, where dark hadrons produced in the Higgs decay promptly decay to SM hadrons. Very recently, AEs have been shown to be effective for detecting dark showers or SUEPs [32, 33].
@@
Robustness of autoencoders: The robustness of an AE as an anomaly tagger is tested in an interesting manner in Ref. [34], which attempts to distinguish boosted top jets from QCD jets with two simple convolutional AEs working on jet images. They train the first AE on a pure QCD-jets sample (direct tagger), and the second one on a pure top-jets sample (inverse tagger). The inverse tagger is used to tag QCD jets as anomalies in a background sample of top jets. The direct tagger AE is found to be a powerful top tagger, agreeing with the findings in Refs. [35, 28]. The authors find that the AE shows a strong complexity bias, i.e., the images that would intuitively be labelled as simpler are reconstructed better. The simpler images are less probable to be identified as anomalous, as AE cannot reconstruct the complex structure of top images when ignoring all but a few pixels. Limitations to the reconstruction of structures in high-energy physics have also been noted recently in Refs. [36, 37]. This correlation between the complexity of an image and the reconstruction loss makes AE unreliable as a model-independent anomaly tagger, as this leads to the assumption that the BSM signal should look more complex than the SM background. They discuss two modifications to the AE to learn more relevant features within the jet images: 1) an intensity remapping for the jet images and 2) the kernel MSE loss function. They find improvement in their results both in terms of the AE performance and the model independence of the tagger. Other papers like Refs. [38, 16] also discuss the challenges of using AE for anomaly detection.
@@
Challenges from smearing: Reconstruction of the final-state objects is one of the main sources of experimental uncertainties (smearing) affecting data-driven methods. The AE learns the kinematic features of the background samples and identifies the signal using the loss function. Smearing shifts the background distribution, making anomaly detection difficult in the case where the smeared background distribution comes closer to the signal. The traditional AE can be combined with an adversarial network (AN, more about it in Sec. 4) to improve the reliability and robustness of this approach, making the autoencoder’s reconstruction independent of the smearing of the background, as shown in Ref. [39]. The authors consider the $3$-momenta of the jets, leptons, and the missing energy to be smeared with Gaussian functions. They train their network with simulated (or pseudo) data with the background events smeared in three possible ways (upwards, downwards, and zero smearing for all objects). The AN takes the AE loss as the input and attempts to discriminate the smearing class of the background. This, in turn, forces the classifier to make its prediction independent of the smearing, making it hard for the AN to discriminate the background samples. Since AE training is generally insensitive to minor signal contaminations, even though their analysis is done on pseudo-data, the authors argue that this procedure can be applied analogously to experimental data by creating labelled datasets that have been systematically smeared.
3.2 Variational autoencoders

### v3

# Unsupervised and lightly supervised learning in particle physics

@@
Abstract
We review the main applications of machine learning models that are not fully supervised in particle physics, i.e., clustering, anomaly detection, detector simulation, and unfolding. Unsupervised methods are ideal for anomaly detection tasks—machine learning models can be trained on background data to identify deviations if we model the background data precisely. The learning can also be partially unsupervised when we can provide some information about the anomalies at the data level. Generative models are useful in speeding up detector simulations—they can mimic the computationally intensive task without large resources. They can also efficiently map detector-level data to parton-level data (i.e., data unfolding). In this review, we focus on interesting ideas and connections and briefly overview the underlying techniques wherever necessary.
I Introduction
In recent years—especially since the advent of deep learning—machine learning (ML) techniques have undergone what can be termed dramatic revolutions. Their performance and range of applicability have increased manyfold. ML tools are now standard in almost every field of study. Particle physics is no exception. These techniques are not new to particle physics—from the LEP to the LHC, high-energy data have been analysed with multivariate techniques like boosted decision trees, neural networks (NNs), etc. (see the references in, e.g., Ref. [1]). The famous discovery of the Higgs boson, which completed the experimental verification of the particle spectrum of the Standard Model (SM), relied on ML techniques. Recently, however, there has been a surge in interest in these methods. Modern and sophisticated ML techniques are being adapted to particle physics problems at an unprecedented rate (see Ref. [2] [^2] The Particle Data Book  [3] now includes a review on ML.) for various reasons. For example, precise measurements in some unexplored parts of the SM, like the Higgs self-coupling or light-quark Yukawa couplings, will be beyond the reach of the traditional analyses [4, 5, 6, 7]. Also, the fact that the LHC is yet to find any signature of physics beyond the SM (BSM) implies that we need to look for processes with even smaller cross-sections or leading to minor departures from known distributions. With the planned high luminosity upgrade (HL-LHC), the LHC will gather, analyse, and record a vast amount of collision data. Modern ML techniques are ideal for such tasks—they can handle tiny signals much better than the usual cut-based analyses designed by physicists’ intuitions or find (learn) patterns from a vast amount of data.
@@ I Introduction
Unsupervised tasks are not new to particle physics. Common jet clustering is essentially an unsupervised task as the algorithm (like anti-$k_{T}$ or Cambridge/Aachen) groups objects without any specific predefined target jet. [^3] In practice, some tuning of the parameters may be necessary to tag boosted objects; for example, see Refs. [8, 9], where a dijet signal from a boosted BSM singlet boson is tagged. We look at clustering tasks in Sec. II. In Sec. III, we focus on anomaly detection, another major task for nonsupervised methods. Anomaly detection without supervision is essentially finding deviations or departures (anomalies) from known patterns without any prior expectation about the nature of the anomalies (like knowledge of any NP model or interaction). If the ML model learns the distribution of the background (SM) events directly from the data, it can identify events not coming from that distribution (i.e., out-of-distribution events). In this section, we discuss data-driven background estimation, autoencoders, variational autoencoders, weakly supervised methods, topic modelling in jet space, and self-supervised learning methods. In the next section (Sec. IV), we look at how ML models, especially generative adversarial networks (GANs), can play a key role in fast detector simulations and unfolding the data distributions recorded at the detectors. Finally, we conclude in Sec. V. One comment is in order before we move on to the next section. Most developments we discuss here are recent, and applications of nonsupervised techniques in particle physics are still nascent. [^4] The HEP community has put out several challenges related to unsupervised learning in the last five years: Fast Calorimeter Challenge: https://calochallenge.github.io/homepage/, Anomaly Detection Challenge: https://mpp-hep.github.io/ADC2021/, etc. Hence, in this review, we focus more on the methods rather than quantitative comparisons of the results. For specialised reviews on anomaly detection and generative modelling in detector simulations, see Ref. [10] and Ref. [11], respectively.
II Clustering
@@
$\mathcal{T}$$\equiv$ $\{\boldsymbol{x}_{i}\}_{i=1}^{N_{T}}\stackrel{{\scriptstyle\textrm{iid}}}{{\sim}}\rho_{T}\,,$ (4)
$\mathcal{B}$$\equiv$ $\{\boldsymbol{x}^{\prime}_{i}\}_{i=1}^{N_{B}}\stackrel{{\scriptstyle\textrm{iid}}}{{\sim}}\rho_{B}\,.$ (5)
with $\{\boldsymbol{x}_{i}|\boldsymbol{x}_{i}\in\mathbb{R}^{D}\}_{i=1}^{N_{T}}$ and $\{\boldsymbol{x}^{\prime}_{i}|\boldsymbol{x}^{\prime}_{i}\in\mathbb{R}^{D}\}_{i=1}^{N_{B}}$ being two independent and identically distributed $D$-dimensional samples drawn independently from the probability density functions (PDFs) $\rho_{T}$ and $\rho_{B}$, respectively. The $\mathcal{T}$, $\mathcal{B}$ samples consist of $N_{T}$, $N_{B}$ points, respectively. In conventional approaches, the NP signals are simulated in specific models and compared with the experimental data to test for those models. Since the theoretical landscape for NP signals is enormous, there is always a possibility that an actual signal may be missed. This method looks for deviations between the simulated SM background ($\mathcal{B}$ samples) and the true data ($\mathcal{T}$), agnostic to any NP scenarios. They check to what significance the two samples are compatible with each other (i.e., $\rho_{B}=\rho_{T}$) by performing a statistical test of the null hypothesis $\{H_{0}:\rho_{T}=\rho_{B}\}$ versus the alternative hypothesis $\{H_{1}:\rho_{T}\neq\rho_{B}\}$. The test statistic (TS) quantifies how much the sample data deviates from what is expected under the null hypothesis. The TS over the trial sample is defined as:
$\textrm{TS}(\mathcal{B},\mathcal{T})\equiv\log\hat{\lambda}^{1/|\mathcal{T}|}=\frac{1}{N_{T}}\sum_{j=1}^{N_{T}}\log\frac{\hat{\rho}_{T}(\boldsymbol{x}_{j})}{\hat{\rho}_{B}(\boldsymbol{x}_{j})}\,,$ (6)
@@
$\hat{\rho}_{B}(\boldsymbol{x}_{j})$ $=$ $\frac{\overline{K}}{N_{B}}\frac{1}{\omega_{D}r_{j,B}^{D}},\quad\quad\hat{\rho}_{T}(\boldsymbol{x}_{j})=\frac{\overline{K}}{N_{T}-1}\frac{1}{\omega_{D}r_{j,T}^{D}}\,,$ (7)
(for any $\boldsymbol{x}_{j}\in\mathcal{T}$) where $\omega_{D}=\pi^{D/2}/\Gamma(D/2+1)$ is the volume of the unit sphere in $\mathbb{R}^{D}$, and $\overline{K}$ is the number of nearest neighbours.
@@ III Anomaly detection
Figure 1: ABCD method: Figure adapted from Ref. [15]. Signal (background) points are coloured blue (light brown). The independence of $f$ and $g$ implies Pr$(f\geq f_{c}{\rm~{}and~{}}g\geq g_{c})={\rm Pr}(f\geq f_{c})\times{\rm Pr}(g\geq g_{c})$.
@@
((a))   (a)
((b))   (b)
Figure 3: Left Panel: Illustration of the AE architecture (taken from Ref. [24]). An AE comprises two neural networks – the encoder and the decoder. The encoder learns a transformation from the input $x$ to a compressed (bottleneck, shown in orange) latent space representation $z$, and the decoder learns a transformation from the latent space representation $z$ to reconstructed output $x^{\prime}$. The AE is trained such that the reconstructed output is ideally identical to the input ($x\approx x^{\prime}$). AEs learn a data-biased algorithm that is excellent at compressing and reconstructing data. The reconstruction quality of $x^{\prime}$ with respect to $x$ can be modelled as an anomaly score. Right Panel: The difference between the reconstruction loss of the signal and the background. Since AEs are trained to construct the background very well, the reconstruction error for the background is lower than that for the signal. A cut on the loss can be used to screen for anomalies.
III.1 Autoencoders
AEs are generally used for dimensionality reduction and feature learning, but they can also be used as unsupervised classifiers in some scenarios. In its basic form, an AE has a symmetrical architecture and consists of an encoder and a decoder, as shown in Fig. 3(a). The encoder consists of one or more layers of neurons that map the input data ($x\in\mathbb{R}^{n}$) into a lower-dimensional representation (latent space, $y\in\mathbb{R}^{k}$ with $k<n$). Most salient features of the input data captured by this compressed representation can be implemented by the function $f_{\theta}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{k}$. The number of neurons in the encoder decreases with each layer until it reaches the bottleneck layer. The bottleneck layer (also known as the latent space or the coding layer) has the lowest number of neurons and represents the space with the compressed form of the input data. After that, the decoder reconstructs the input data from the compressed representation, $g_{\varphi}:\mathbb{R}^{k}\rightarrow\mathbb{R}^{n}$. The decoder consists of the same number of layers as the encoder, with the number of neurons in each decoder layer increasing until it reaches the output layer, which has the same number of neurons as the input layer. Each hidden layer in this network typically uses activation functions like the rectified linear unit (ReLU) to introduce non-linearity. The autoencoder training involves minimising a loss function [e.g., mean-squared error (MSE)] that measures the difference between the input data and the reconstructed output:
@@ III.1 Autoencoders
During training, the network learns to map the input data to a lower-dimensional representation and later reconstructs the input from that representation. As mentioned earlier, AEs can be used to identify events inconsistent with the expected background distribution of the SM as a sign of physics beyond the SM. The signal events would result in larger $L$ (Eq. (13)) than usual, i.e., a signal is more likely to manifest in the tail of the $L$ distribution, as shown in Fig. 3(b).
Figure 4: The distribution of the anomaly score from the AE is presented for the data and the five benchmark models ($H^{+}$, $W_{\text{KK}}$, $Z^{\prime}$, $W^{\prime}$ and $Z^{\prime}$, respectively). The predictions from these BSM models describe the expected number of events from 140fb-1 of data for heavy particle masses around 2 TeV. The vertical lines represent the start of the three anomaly regions (AR). The labels associated with the three ARs state the visible cross section for the BSM processes which yields the same number of events as observed in the 140fb-1 dataset. For further details refer to Ref. [25].
@@
Graphs (a set of objects and pair-wise relations among them) can be used to represent data in an intuitive, permutation-invariant way. Graph representations are well-suited to problems in particle physics. For example, parton showering can be visualised as a directed graph along the direction of increasing quark/gluon multiplicity. One could also construct a graph of an event from reconstructed final states and the relationships among them. Graph neural networks (GNNs) are a class of DL models that learn functions from graphical data. Typically training GNNs involves several iterations of message-passing (where nodes exchange information with their neighbours) and learning functions that update each node’s state after every iteration (based on its current status and the incoming messages). For a detailed review of general GNN architecture and applications in particle physics, see Ref. [30]. The prediction task for a GNN can be at the node (or vertex) level, edge level, or graph level. Ref. [31] uses a GNN-based AE for anomaly detection at the graph level. The AE is trained on QCD jets to tag hadronically decaying $W$ jets, top jets, and a boosted scalar (which decays to two $W$ jets to give a four-pronged signature) as anomalies. The authors introduce a novel edge-reconstruction decoder that helps the AE reconstruct the jet graph of the jet. Ref. [32] shows that graph AEs can be made IRC-safe by constructing the graphs in an IRC-safe manner and by constraining the learned function to rely on energy weights for node/edge relationships within the message-passing algorithm.
@@
Hidden-valley models can have an additional strong (confining) gauge group under which all SM particles are singlets [33]. However, these hidden dynamics might form light (dark) hadrons, which could be produced at the LHC through some higher-dimensional operators or an intermediary like the dark photon that mixes with the ordinary photon or the Higgs [e.g., the Higgs decaying to dark hadrons], etc. Like the bound states in QCD, dark hadrons can produce dark showers. Since modelling the signals is difficult in these cases, they make a good case for unsupervised learning in general. Ref. [34] demonstrated this by looking for soft-unclustered energy patterns (SUEPs) produced by dark showers, which manifest as isotropic and soft energy deposits. The authors analysed one of the experimentally challenging scenarios, where dark hadrons produced in the Higgs decay promptly decay to SM hadrons. Very recently, AEs have been shown to be effective for detecting dark showers or SUEPs [35, 36].
@@
Robustness of autoencoders: The robustness of an AE as an anomaly tagger is tested in an interesting manner in Ref. [37], which attempts to distinguish boosted top jets from QCD jets with two simple convolutional AEs working on jet images. They train the first AE on a pure QCD-jets sample (direct tagger), and the second one on a pure top-jets sample (inverse tagger). The inverse tagger is used to tag QCD jets as anomalies in a background sample of top jets. The direct tagger AE is found to be a powerful top tagger, agreeing with the findings in Refs. [38, 28]. The inverse tagger fails to tag an anomaly, i.e., the QCD jets, and performs worse than randomly tagging jets as anomalous. The authors find that the AE shows a strong complexity bias, i.e., the images that would intuitively be labelled as simpler are reconstructed better. The simpler images are less probable to be identified as anomalous, as AE cannot reconstruct the complex structure of top images when ignoring all but a few pixels. Limitations to the reconstruction of structures in high-energy physics have also been noted recently in Refs. [39, 40]. This correlation between the complexity of an image and the reconstruction loss makes AE unreliable as a model-independent anomaly tagger, as this leads to the assumption that the BSM signal should look more complex than the SM background. They discuss two modifications to the AE to learn more relevant features within the jet images: 1) an intensity remapping for the jet images and 2) the kernel MSE loss function. They find improvement in their results both in terms of the AE performance and the model independence of the tagger. Other papers like Refs. [41, 16] also discuss the challenges of using AE for anomaly detection.
@@
Challenges from smearing: Reconstruction of the final-state objects is one of the main sources of experimental uncertainties (smearing) affecting data-driven methods. The AE learns the kinematic features of the background samples and identifies the signal using the loss function. Smearing shifts the background distribution, making anomaly detection difficult in the case where the smeared background distribution comes closer to the signal. The traditional AE can be combined with an adversarial network (AN, more about it in Sec. IV) to improve the reliability and robustness of this approach, making the autoencoder’s reconstruction independent of the smearing of the background, as shown in Ref. [42]. The authors consider the $3$-momenta of the jets, leptons, and the missing transverse energy to be smeared with Gaussian functions. They train their network with simulated (or pseudo) data with the background events smeared in three possible ways (upwards, downwards, and zero smearing for all objects). The AN takes the AE loss as the input and attempts to discriminate the smearing class of the background. This, in turn, forces the classifier to make its prediction independent of the smearing, making it hard for the AN to discriminate the background samples. Since AE training is generally insensitive to minor signal contaminations, even though their analysis is done on pseudo-data, the authors argue that this procedure can be applied analogously to experimental data by creating labelled datasets that have been systematically smeared.
III.2 Variational autoencoders